# Fair Classifiers Without Fair Training: An Influence-Guided Data Sampling Approach

### v1

@@ Appendix A More Details of Related Work
Fair classification. The fairness-aware learning algorithms, in general, can be categorized into pre-processing, in-processing, and post-processing methods. Pre-processing methods typically reweigh or distort the data examples to mitigate the identified biases [Asudeh et al., 2019, Azzalini et al., 2021, Tae and Whang, 2021, Sharma et al., 2020, Celis et al., 2020, Chawla et al., 2002, Zemel et al., 2013, Chen et al., 2018]. More relevant to us is the importance reweighting, which assigns different weights to different training examples to ensure fairness across groups [Kamiran and Calders, 2012, Jiang and Nachum, 2020, Diesendruck et al., 2020, Choi et al., 2020, Qraitem et al., 2023, Li and Vasconcelos, 2019]. Our proposed algorithm bears similarity to a specific case of importance reweighting, particularly the 0-1 reweighting applied to newly added data. The main advantage of our work, however, lies in its ability to operate without needing access to the sensitive attributes of either the new or training data. Other parallel studies utilize importance weighting to learn a complex fair generative model in a weakly supervised setting [Diesendruck et al., 2020, Choi et al., 2020], or to mitigate representation bias in training datasets [Li and Vasconcelos, 2019].
Post-processing methods typically enforce fairness on a learned model through calibration [Feldman, 2015, Feldman et al., 2015, Hardt et al., 2016]. Although this approach is likely to decrease the disparity of the classifier, by decoupling the training from the fairness enforcement, this procedure may not lead to the best trade-off between fairness and accuracy [Woodworth et al., 2017, Pleiss et al., 2017]. In contrast, our work can achieve a better trade-off between fairness and accuracy, because we reduce the fairness disparity by mitigating the adverse effects of distribution shifts on generalization error. Additionally, these post-processing techniques necessitate access to the sensitive attribute during the inference phase, which is often not available in many real-world scenarios.
Fairness-accuracy tradeoff. It has been demonstrated that there is an implicit trade-off between fairness and accuracy in the literature. Compared to the prior works [Menon and Williamson, 2018, Prost et al., 2019], our work does not require additional assumptions about the classifier and the characteristics of the training/testing datasets (for example, distribution shifts). Besides, the main difference is that this paper does not actually work in the classical regime of the fairness-accuracy tradeoff. By properly collecting new data from an unlabeled dataset, we can improve both generalization and fairness at the same time, which cannot be achieved by working on a fixed and static training dataset that naturally incurs such a tradeoff.
Li and Liu [2022] is a similar work that utilizes the influence function to reweight the data examples but requires re-training. Our method focuses on soliciting additional samples from an external unlabeled dataset while [Li and Liu, 2022] reweights the existing and fixed training dataset. Our approach is more closely with a fair active learning approach [Anahideh et al., 2022]. However, this fair active learning framework relies on sensitive attribute information while our algorithm does not.
Distribution shifts. Common research concerning distribution shifts necessitates extra assumptions to build theoretical connections between features and attributes, like causal graphs [Singh et al., 2021], correlation shifts [Roh et al., 2023], and demographic shifts [Giguere et al., 2022]. In contrast, our approach refrains from making further assumptions about the properties of distribution shifts.
In this literature, many works have utilized distributionally robust optimization (DRO) to reduce fairness disparity without sensitive attribute information [Hashimoto et al., 2018, Kirichenko et al., 2022, Liu et al., 2021b, Lahoti et al., 2020, Veldanda et al., 2023, Sohoni et al., 2020, Wei et al., 2023b]. Although these works also evaluate the worst-group performance in the context of fairness, their approach differs as they do not strive to equalize the loss across all groups. Besides, in these studies, accuracy and worst-case accuracy are used to showcase the efficacy of DRO. Essentially, they equate fairness with uniform accuracy across groups, implying a model is considered fair if it demonstrates equal accuracies for all groups. However, this specific definition of fairness is somewhat restrictive and does not align with more conventional fairness definitions like DP or EOd.
Appendix B Omitted Proofs
@@
In this section, we present detailed proofs for the lemmas and theorems in Sections 4 and 5, respectively.
B.1 Proof of Lemma 4.1
Lemma B.1.
[…]
@@
$\begin{split}\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})&=\underbrace{\bigg{(}\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{P}}(\mathbf{w}^{P})\bigg{)}}_{\text{distribution shift}}+\underbrace{\bigg{(}\mathcal{R}_{\mathcal{P}}(\mathbf{w}^{P})-\mathcal{R}_{P}(\mathbf{w}^{P})\bigg{)}}_{\text{Hoeffding's inequality}}+\underbrace{\mathcal{R}_{P}(\mathbf{w}^{P}))}_{\text{empirical risk}}\\ &\leq G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})+\sqrt{\frac{\log(4/\delta)}{2N_{P}}}+\mathcal{R}_{P}(\mathbf{w}^{P})\\ \end{split}$
@@
$\begin{split}\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{P}}(\mathbf{w}^{P})&=\mathbb{E}_{z\sim\mathcal{Q}}[\ell(\mathbf{w}^{P},z)]-\mathbb{E}_{z\sim\mathcal{P}}[\ell(\mathbf{w}^{P},z)]\\ &=\sum_{i=1}^{I}p^{(\mathcal{Q})}(\pi=i)\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]-\sum_{i=1}^{I}p^{(\mathcal{P})}(\pi=i)\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]\\ &=\sum_{i=1}^{I}|p^{(\mathcal{P})}(\pi=i)-p^{(\mathcal{Q})}(\pi=i)|\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]\\ &\leq G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q}).\end{split}$
where we define $\text{dist}(\mathcal{P},\mathcal{Q})=\sum_{i=1}^{I}|p^{(\mathcal{P})}(\pi=i)-p^{(\mathcal{Q})}(\pi=i)|$ and $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]\leq G_{P},\forall i\in I$ because of Assumption 4.2. To avoid misunderstanding, we use a subscript $P$ of the constant $G$ to clarify the corresponding model $\mathbf{w}^{P}$. Then, for the second term (Hoeffding inequality), with probability at least $1-\delta$, we have $|\mathcal{R}_{\mathcal{P}}(\mathbf{w}^{P})-\mathcal{R}_{P}(\mathbf{w}^{P})|\leq\sqrt{\frac{\log(4/\delta)}{2N_{P}}}$. ∎
B.2 Proof of Theorem 4.2
Theorem 4.2. (Upper bound of fairness disparity). Suppose $\mathcal{R}_{\mathcal{Q}}(\cdot)$ follows Assumption 4.1. Let $\text{dist}(\mathcal{P},\mathcal{Q})$, $G_{P}$, $\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})$ and $\text{dist}(P_{k},P)$ be defined therein. Given model $\mathbf{w}^{P}$ and $\mathbf{w}^{k}$ trained exclusively on group $k$’s data $P_{k}$, with probability at least $1-\delta$ with $\delta\in(0,1)$, then the upper bound of the fairness disparity is
$\begin{split}\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})\leq\underbrace{G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})}_{\text{ distribution shift}}+\underbrace{\Phi\cdot\text{dist}(P_{k},P)^{2}}_{\text{group gap}}+G_{k}\cdot\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})+\Upsilon.\end{split}$
where
$\Phi=4L^{2}G^{2}\sum_{t=0}^{T-1}(\eta_{t}^{2}(1+2\eta_{t}^{2}L^{2}))^{t},\quad\Upsilon=\sqrt{\frac{\log(4/\delta)}{2N_{P}}}+\sqrt{\frac{\log(4/\delta)}{2N_{P_{k}}}}+\varpi+\varpi_{k}$
Note that $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\leq G_{k}$, $\varpi=\mathcal{R}_{P}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w}^{\mathcal{Q}})$ and $\varpi_{k}=\mathcal{R}_{P_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})$. Specifically, $\varpi$ and $\varpi_{k}$ can be regarded as constants because $\mathcal{R}_{P}(\mathbf{w}^{P})$ and $\mathcal{R}_{P_{k}}(\mathbf{w}^{k})$ correspond to the empirical risks, $\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})$ and $\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w^{\mathcal{Q}_{k}}})$ represent the ideal minimal empirical risk of model $\mathbf{w}^{\mathcal{Q}}$ trained on distribution $\mathcal{Q}$ and $\mathcal{Q}_{k}$, respectively.
@@
$\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})$ $=(\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P_{k}})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P}))+(\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P_{k}}))$
@@
where $\mathbf{w}^{P_{k}}$ represents the model trained exclusively on group $k$’s data. For simplicity, when there is no confusion, we use $\mathbf{w}^{k}$ to substitute $\mathbf{w}^{P_{k}}$. The inequality $\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{k})\leq 0$ holds due to the fact that the model tailored for a single group $k$ can not generalize well to the entirety of the test set $Q$.

### v3

@@ Appendix A More details of related work
Fair classification The fairness-aware learning algorithms, in general, can be categorized into pre-processing, in-processing, and post-processing methods. Pre-processing methods typically reweigh or distort the data examples to mitigate the identified biases [7, 9, 65, 60, 15, 17, 80, 18]. More relevant to us is the importance reweighting, which assigns different weights to different training examples to ensure fairness across groups [37, 36, 23, 21, 57, 44]. Our proposed algorithm bears similarity to a specific case of importance reweighting, particularly the 0-1 reweighting applied to newly added data. The main advantage of our work, however, lies in its ability to operate without needing access to the sensitive attributes of either the new or training data. Other parallel studies utilize importance weighting to learn a complex fair generative model in a weakly supervised setting [23, 21], or to mitigate representation bias in training datasets [44]. Post-processing methods typically enforce fairness on a learned model through calibration [26, 27, 32]. Although this approach is likely to decrease the disparity of the classifier, by decoupling the training from the fairness enforcement, this procedure may not lead to the best trade-off between fairness and accuracy [75, 55]. In contrast, our work can achieve a better trade-off between fairness and accuracy, because we reduce the fairness disparity by mitigating the adverse effects of distribution shifts on generalization error. Additionally, these post-processing techniques necessitate access to the sensitive attribute during the inference phase, which is often not available in many real-world scenarios.
Appendix B Understanding risk disparity through common fairness metrics
Here, we provide a brief proof to illustrate Proposition 3.1, that is, the relationship between risk parity and common fairness metrics, such as DP and EOd. For completeness, here we provide a more detailed version of Proposition 3.1, as well as its proof.
Proposition 3.1. Consider a binary classification scenario involving two demographic groups $S\in\{k,k^{\prime}\}$. When two groups are balanced, i.e., $\frac{\mathbb{P}(Y=y|S=k)}{\mathbb{P}(Y=y|S=k^{\prime})}=1$, where $\hat{Y}$ denotes the predicted label, risk disparity can serve as a lower bound for fairness disparities based on EOd. Similarly, if the group sufficiency ratio can be calibrated by 1, i.e., $\mathbb{P}(Y=y|S=k,\hat{Y}=y)=\mathbb{P}(Y=y|S=k^{\prime},\hat{Y}=y)=R$ where $R$ is calibration score, risk disparity can be formulated as a lower bound for DP-based fairness disparity.
Proof.
Consider a scenario involving two demographic groups $S\in\{k,k^{\prime}\}$, alongside a 0-1 loss function.
Fairness metric EOd
For EOd, the risk disparity can be reformulated as
$|\mathcal{R}_{Q_{k}}(\mathbf{w}^{P})-\mathcal{R}_{Q_{k^{\prime}}}(\mathbf{w}^{P})|$
$=$ $\left|\frac{1}{|Q_{k}|}\sum\nolimits_{\left(x_{n},y_{n}\right)\in Q_{k}}\mathbb{I}\left(\hat{y}_{n}\neq y_{n}\right)-\frac{1}{|Q_{k^{\prime}}|}\sum\nolimits_{\left(x_{n},y_{n}\right)\in Q_{k^{\prime}}}\mathbb{I}\left(\hat{y}_{n}\neq y_{n}\right)\right|$
$=$ $\left|\mathbb{P}(\hat{Y}\neq Y\mid S=k)-\mathbb{P}(\hat{Y}\neq Y\mid S=k^{\prime})\right|$
$=$ $\left|\sum_{y}\left(\mathbb{P}\left(Y=y\mid S=k^{\prime}\right)\mathbb{P}(\hat{Y}=y\mid S=k^{\prime},Y=y)-\mathbb{P}(Y=y\mid S=k)\mathbb{P}(\hat{Y}=y\mid S=k,Y=y)\right)\right|$
$=$ $\left|\sum_{y}\mathbb{P}(Y=y\mid S=k^{\prime})\cdot\left(\mathbb{P}(\hat{Y}=y\mid S=k^{\prime},Y=y)-\omega_{\text{EOd}}(y)\mathbb{P}(\hat{Y}=y\mid S=k,Y=y)\right)\right|$
$=$ $\bigg{|}\mathbb{P}(Y=1|S=k^{\prime})\cdot\left(\mathbb{P}(\hat{Y}=1\mid S=k^{\prime},Y=1)-\omega_{\text{EOd}}(y=1)\cdot\mathbb{P}(\hat{Y}=1\mid S=k,Y=1)\right)\quad\quad\text{\# binary case}$
$\quad+\mathbb{P}(Y=0|S=k^{\prime})\cdot\left(\mathbb{P}(\hat{Y}=0\mid S=k^{\prime},Y=0)-\omega_{\text{EOd}}(y=0)\cdot\mathbb{P}(\hat{Y}=0\mid S=k,Y=0)\right)\bigg{|}$
$=$ $\bigg{|}\mathbb{P}(Y=1|S=k^{\prime})\cdot\left(\mathbb{P}(\hat{Y}=1\mid S=k^{\prime},Y=1)-\omega_{\text{EOd}}(y=1)\cdot\mathbb{P}(\hat{Y}=1\mid S=k,Y=1)\right)\quad\quad$
$\quad+\mathbb{P}(Y=0|S=k^{\prime})\cdot\left(\omega_{\text{EOd}}(y=0)\cdot\mathbb{P}(\hat{Y}=1\mid S=k,Y=0)-\mathbb{P}(\hat{Y}=1\mid S=k^{\prime},Y=0)\right)$
$\quad+\mathbb{P}(Y=0|S=k^{\prime})\cdot(1-\omega_{\text{EOd}}(y=0))\bigg{|}$
$=$ $\bigg{|}\mathbb{P}(Y=1|S=k^{\prime})\cdot\left(\mathbb{P}(\hat{Y}=1\mid S=k^{\prime},Y=1)-\mathbb{P}(\hat{Y}=1\mid S=k,Y=1)\right)\quad\quad$
$\quad-\mathbb{P}(Y=0|S=k^{\prime})\cdot\left(\mathbb{P}(\hat{Y}=1\mid S=k^{\prime},Y=0)-\mathbb{P}(\hat{Y}=1\mid S=k,Y=0)\right)\bigg{|}\quad\quad\text{\# balanced groups}$
$\leq$ $\mathbb{P}(Y=1|S=k^{\prime})\cdot\bigg{|}\mathbb{P}(\hat{Y}=1\mid S=k^{\prime},Y=1)-\mathbb{P}(\hat{Y}=1\mid S=k,Y=1)\bigg{|}$
$\quad+\mathbb{P}(Y=0|S=k^{\prime})\cdot\bigg{|}\mathbb{P}(\hat{Y}=1\mid S=k^{\prime},Y=0)-\mathbb{P}(\hat{Y}=1\mid S=k,Y=0)\bigg{|}$
$\leq$ $\underbrace{\sum_{y\in\{0,1\}}\bigg{|}\mathbb{P}(\hat{Y}=1\mid S=k^{\prime},Y=y)-\mathbb{P}(\hat{Y}=1\mid S=k,Y=y)\bigg{|}}_{\text{ EOd-based fairness disparity}}$
where we define $\omega_{\text{EOd}}(y):=\frac{\mathbb{P}(Y=y|S=k)}{\mathbb{P}(Y=y|S=k^{\prime})}$, serving as the bias weight. Here, we make a mild assumption that the two demographic groups are balanced, i.e., $\omega_{\text{EOd}}(y)=1$. Then, we can see that the last item measures the fairness disparity based on EOd for binary classification problems. Thus, we can claim that reducing risk disparity can promote the fairness metric EOd.
Fairness metric DP
Similarly, for DP, we formulate the risk disparity as
$|\mathcal{R}_{Q_{k}}(\mathbf{w}^{P})-\mathcal{R}_{Q_{k^{\prime}}}(\mathbf{w}^{P})|$
$=$ $\left|\frac{1}{\left|Q_{k}\right|}\sum\nolimits_{\left(x_{n},y_{n}\right)\in Q_{k}}\mathbb{I}\left(\hat{y}_{n}\neq y_{n}\right)-\frac{1}{|Q_{k^{\prime}}|}\sum\nolimits_{\left(x_{n},y_{n}\right)\in Q_{k^{\prime}}}\mathbb{I}\left(\hat{y}_{n}\neq y_{n}\right)\right|$
$=$ $\left|\mathbb{P}(\hat{Y}\neq Y\mid S=k)-\mathbb{P}(\hat{Y}\neq Y\mid S=k^{\prime})\right|$
$=$ $\left|\sum_{y}\left(\mathbb{P}(\hat{Y}=y\mid S=k^{\prime})\mathbb{P}(Y=y\mid S=k^{\prime},\hat{Y}=y)-\mathbb{P}(\hat{Y}=y\mid S=k)\mathbb{P}(Y=y\mid S=k,\hat{Y}=y)\right)\right|$
$=$ $\left|\sum_{y}\mathbb{P}(Y=y|S=k^{\prime},\hat{Y}=y)\cdot\left(\mathbb{P}(\hat{Y}=y\mid S=k^{\prime})-\omega_{\text{DP}}(y)\mathbb{P}(\hat{Y}=y\mid S=k)\right)\right|$
$=$ $\bigg{|}\mathbb{P}(Y=1|S=k^{\prime},\hat{Y}=1)\cdot\left(\mathbb{P}(\hat{Y}=1\mid S=k^{\prime})-\omega_{\text{DP}}(y=1)\cdot\mathbb{P}(\hat{Y}=1\mid S=k)\right)\quad\quad\text{\# binary case}$
$\quad+\mathbb{P}(Y=0|S=k^{\prime},\hat{Y}=0)\cdot\left(\mathbb{P}(\hat{Y}=0\mid S=k^{\prime})-\omega_{\text{DP}}(y=0)\cdot\mathbb{P}(\hat{Y}=0\mid S=k)\right)\bigg{|}$
$=$ $\bigg{|}\mathbb{P}(Y=1|S=k^{\prime},\hat{Y}=1)\cdot\left(\mathbb{P}(\hat{Y}=1\mid S=k^{\prime})-\omega_{\text{DP}}(y=1)\cdot\mathbb{P}(\hat{Y}=1\mid S=k)\right)\quad\quad$
$\quad+\mathbb{P}(Y=0|S=k^{\prime},\hat{Y}=0)\cdot\left(\omega_{\text{DP}}(y=0)\cdot\mathbb{P}(\hat{Y}=1\mid S=k)-\mathbb{P}(\hat{Y}=1\mid S=k^{\prime})\right)$
$\quad+\mathbb{P}(Y=0|S=k^{\prime},\hat{Y}=0)\cdot(1-\omega_{\text{DP}}(y=0))\bigg{|}\quad\quad\quad\quad\text{\# Calibrating group sufficiency}$
$=$ $\bigg{|}\mathbb{P}(Y=1|S=k^{\prime},\hat{Y}=1)\cdot\left(\mathbb{P}(\hat{Y}=1\mid S=k^{\prime})-\mathbb{P}(\hat{Y}=1\mid S=k)\right)\quad\quad$
$\quad+\mathbb{P}(Y=0|S=k^{\prime},\hat{Y}=0)\cdot\left(\mathbb{P}(\hat{Y}=1\mid S=k)-\mathbb{P}(\hat{Y}=1\mid S=k^{\prime})\right)\bigg{|}$
$=$ $\bigg{|}\mathbb{P}(Y=1|S=k^{\prime},\hat{Y}=1)-\mathbb{P}(Y=0|S=k^{\prime},\hat{Y}=0)\bigg{|}\cdot\bigg{|}\mathbb{P}(\hat{Y}=1\mid S=k^{\prime})-\mathbb{P}(\hat{Y}=1\mid S=k)\bigg{|}$
$\leq$ $\underbrace{\bigg{|}\mathbb{P}(\hat{Y}=1\mid S=k^{\prime})-\mathbb{P}(\hat{Y}=1\mid S=k)\bigg{|}}_{\text{DP-based fairness disparity}}$
where we define $\omega_{\text{DP}}(y):=\frac{\mathbb{P}(Y=y|S=k,\hat{Y}=y)}{\mathbb{P}(Y=y|S=k^{\prime},\hat{Y}=y)}$. In fact, $\omega_{\text{DP}}(y)$ measures the group sufficiency ratio [61]. Note that the group sufficiency is closely related to the idea of calibration [11]. Thus, we make a mild assumption that the group sufficiency ratio $\omega_{\text{DP}}(y)$ can be calibrated by 1, i.e., $\mathbb{P}(Y=y|S=k,\hat{Y}=y)=\mathbb{P}(Y=y|S=k^{\prime},\hat{Y}=y)=R$, where $R$ is a calibration score. When two demographic groups are balanced, i.e., $\omega_{\text{DP}}(y)=1$, we can see that the last inequality indicates the DP-based fairness disparity for binary classification problems. Therefore, we can also claim that the fairness metric DP can also be encouraged when reducing the risk disparity.
Appendix C Detailed analysis of the fair influential sampling algorithm
C.1 Evaluating first-order influence estimations against real influence
Recall that the influence score derived in this paper primarily utilizes a first-order approach. Here, we will demonstrate how accurate the first-order estimation of the influence is in comparison to the real influence.
Figure 4: We validate how accurate the first-order estimation of the influence is in comparison to the real influence. The $x$-axis represents the actual influence per sample, and the $y$-axis represents the estimated influence. We observe that while some of the examples are away from the diagonal line (which indicates the estimation is inaccurate), the estimated influences for most of the data samples are very close to their actual influence values.
C.2 Comparative analysis of computational costs
Recall that the proposed algorithm FIS needs to pre-calculate the accuracy loss and fairness loss for evaluating the performance of a certain example. However, the extra computation cost is comparable to the cost of traditional model training. Note that the main extra computation cost in FIS (Algorithm 1) mainly results from model gradients. Let $p$ denote the number of model parameters, then the cost for computing the gradients is $O(p)$ per sample. Specifically, in each round that involves sampling, we need to calculate three parts of gradients: the gradients of $|U|$ unlabeled instances, the average gradient of $|Q_{v}|$ validation instances w.r.t. accuracy loss, and the gradient of $|Q_{v}|$ validation instances w.r.t. fairness loss. Note that in general, $|Q_{v}|\ll|U|$. In practical implementation, to speed up the calculation of gradients over $|U|$ instances, we randomly sample 0.2-0.5% of the unlabeled dataset in each sampling batch. Additionally, we can increase the number of newly selected examples for each round ($r$) to save computation costs. In our experiments, we usually have 10-20 sampling rounds. For instance, running one experiment for the CelebA dataset on a single GPU roughly requires about 4 hours.
C.3 Exploration of the labeling strategies
  1. Strategy I
  2. Strategy II
Remark C.1.
Proof.
Based on the definition of the influence of the prediction component, as delineated in Eq. (2), it becomes evident that the most uncertain points are obtained when the proxy labels closely align with the true labels. Consequently, the model predictions used in Strategy II also approximate the true labels to minimize the cross-entropy loss. Thus, in a certain sense, Strategy I and Strategy II can be considered equivalent. ∎
Appendix D Omitted proofs
@@
In this section, we present complete proofs for the lemmas and theorems in Section 4 and 5, respectively.
D.1 Proof of Lemma 4.1
Lemma 4.1.
The influence of predictions on the validation dataset $Q_{v}$ can be denoted by
$\textsf{Infl}_{\textsf{acc}}(z^{\prime}):=\sum_{n\in|Q_{v}|}\textsf{Infl}_{\textsf{acc}}(z^{\prime},z_{n}^{\circ})\approx-\eta\sum\nolimits_{n\in|Q_{v}|}\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z_{n}^{\circ})\right\rangle$
Proof.
Taking the first-order Taylor expansion, we will have
$f(\mathbf{w}_{t+1},x_{n}^{\circ})-f(\mathbf{w}_{t},x_{n}^{\circ})\approx$ $\left.\left\langle\frac{\partial f(\mathbf{w},x_{n}^{\circ})}{\partial\mathbf{w}},\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\rangle\right|_{\mathbf{w}=\mathbf{w}_{t}}$
where the last equality holds due to Eq. (1). Therefore,
$\textsf{Infl}_{\textsf{acc}}(z^{\prime}):=\sum_{n\in|Q_{v}|}\textsf{Infl}_{\textsf{acc}}(z^{\prime},z_{n}^{\circ})\approx-\eta\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\sum_{n\in|Q_{v}|}\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z_{n}^{\circ})\right\rangle$
D.2 Proof of Lemma 4.2
Lemma 4.2.
The influence of fairness on the validation dataset $Q_{v}$ can be denoted by
$\textsf{Infl}_{\textsf{fair}}(z^{\prime}):=\sum_{n\in|Q_{v}|}\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ})\approx-\eta\sum_{n\in|Q_{v}|}\left\langle\partial_{\mathbf{w}_{t}}\ell(\mathbf{w}_{t},z^{\prime}),\partial_{\mathbf{w}_{t}}\phi(\mathbf{w}_{t},z_{n}^{\circ})\right\rangle$
Proof.
By first-order approximation, we have
Recall by first-order approximation, we have
Therefore,
$\textsf{Infl}_{\textsf{fair}}(z^{\prime})=\sum_{n\in|Q_{v}|}\textsf{Infl}_{\textsf{fair}}(z^{\prime},z_{n}^{\circ})\approx-\left.\eta\sum_{n\in|Q_{v}|}\left\langle\frac{\partial\ell(\mathbf{w},z^{\prime})}{\partial\mathbf{w}},\frac{\partial\phi(\mathbf{w}_{t},z_{n}^{\circ})}{\partial\mathbf{w}}\right\rangle\right|_{\mathbf{w}=\mathbf{w}_{t}}.$
D.3 Proof of Theorem 5.1
Theorem 5.1.
(Generalization error bound). Let $\text{dist}(\mathcal{P},\mathcal{Q})$, $G_{P}$ be defined therein. With probability at least $1-\delta$ with $\delta\in(0,1)$, the generalization error bound of the model trained on dataset $P$ is
$\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})\leq\underbrace{G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})}_{\text{ distribution shift}}+\sqrt{\frac{\log(4/\delta)}{2|P|}}+\mathcal{R}_{P}(\mathbf{w}^{P}).$ (7)
@@
$\begin{split}\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})&=\underbrace{\bigg{(}\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{P}}(\mathbf{w}^{P})\bigg{)}}_{\text{distribution shift}}+\underbrace{\bigg{(}\mathcal{R}_{\mathcal{P}}(\mathbf{w}^{P})-\mathcal{R}_{P}(\mathbf{w}^{P})\bigg{)}}_{\text{Hoeffding's inequality}}+\underbrace{\mathcal{R}_{P}(\mathbf{w}^{P}))}_{\text{empirical risk}}\\ &\leq G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})+\sqrt{\frac{\log(4/\delta)}{2|P|}}+\mathcal{R}_{P}(\mathbf{w}^{P})\\ \end{split}$
@@
where we define $\text{dist}(\mathcal{P},\mathcal{Q})=\sum_{i=1}^{I}|p^{(\mathcal{P})}(\pi=i)-p^{(\mathcal{Q})}(\pi=i)|$ and $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{P},z)]\leq G_{P},\forall i\in I$ because of Assumption 5.2. To avoid misunderstanding, we use a subscript $P$ of the constant $G$ to clarify the corresponding model $\mathbf{w}^{P}$. Then, for the second term (Hoeffding inequality), with probability at least $1-\delta$, we have $|\mathcal{R}_{\mathcal{P}}(\mathbf{w}^{P})-\mathcal{R}_{P}(\mathbf{w}^{P})|\leq\sqrt{\frac{\log(4/\delta)}{2|P|}}$. ∎
D.4 Proof of Theorem 5.2
Theorem 5.2.
(Upper bound of fairness disparity). Suppose $\mathcal{R}_{\mathcal{Q}}(\cdot)$ follows Assumption 5.1. Let $\text{dist}(\mathcal{P},\mathcal{Q})$, $G_{P}$, $\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})$ and $\text{dist}(P_{k},P)$ be defined therein. The initial learning rate $\eta_{0}^{2}<\frac{1}{\sqrt{2}TL}$, where $T$ denotes the number of training epochs. Given model $\mathbf{w}^{P}$ and $\mathbf{w}^{k}$ trained exclusively on group $k$’s data $P_{k}$, with probability at least $1-\delta$ with $\delta\in(0,1)$, then the upper bound of the fairness disparity is
$\begin{split}\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})\leq\underbrace{G_{P}\cdot\text{dist}(\mathcal{P},\mathcal{Q})}_{\text{ distribution shift}}+\underbrace{4L^{2}G^{2}\cdot\text{dist}(P_{k},P)^{2}}_{\text{group gap}}+G_{k}\cdot\text{dist}(\mathcal{P}_{k},\mathcal{Q}_{k})+\Upsilon.\end{split}$
where $\Upsilon=\sqrt{\frac{\log(4/\delta)}{2|P|}}+\sqrt{\frac{\log(4/\delta)}{2|P_{k}|}}+\varpi+\varpi_{k}$. Note that $\mathbb{E}_{z\sim\pi_{i}}[\ell(\mathbf{w}^{k},z)]\leq G_{k}$, $\varpi=\mathcal{R}_{P}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w}^{\mathcal{Q}})$ and $\varpi_{k}=\mathcal{R}_{P_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w}^{\mathcal{Q}_{k}})$. Specifically, $\varpi$ and $\varpi_{k}$ can be regarded as constants because $\mathcal{R}_{P}(\mathbf{w}^{P})$ and $\mathcal{R}_{P_{k}}(\mathbf{w}^{k})$ correspond to the empirical risks, $\mathcal{R}_{\mathcal{Q}}^{*}(\mathbf{w^{\mathcal{Q}}})$ and $\mathcal{R}_{\mathcal{Q}_{k}}^{*}(\mathbf{w^{\mathcal{Q}_{k}}})$ represent the ideal minimal empirical risk of model $\mathbf{w}^{\mathcal{Q}}$ trained on distribution $\mathcal{Q}$ and $\mathcal{Q}_{k}$, respectively. Moreover, these quantities $\varpi$ and $\varpi_{k}$ are not taken into account during the training phase, but rather in relation to the final model.
@@
$\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P})$
$=$ $(\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P_{k}})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P}))+(\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{P})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{P_{k}}))$
@@
where $\mathbf{w}^{P_{k}}$ represents the model trained exclusively on group $k$’s data. For simplicity, when there is no confusion, we use $\mathbf{w}^{k}$ to substitute $\mathbf{w}^{P_{k}}$. The inequality $\mathcal{R}_{\mathcal{Q}_{k}}(\mathbf{w}^{k})-\mathcal{R}_{\mathcal{Q}}(\mathbf{w}^{k})\geq 0$ holds because the model tailored for a single group $k$ can not generalize well to the entirety of the test set $Q$.